40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5313     Entry time: Sat Aug 27 20:38:17 2011
Author: Suresh 
Type: Update 
Category: IOO 
Subject: Light is back on WFS 

[Valera, Suresh]

   We wanted to continue the work with WFS servo loops.  As the current optical paths on the AP table do not send any light to the WFS, I changed a mirror to a 98% window and a window to a mirror to send about 0.25mW of light towards the WFS.   The MC locking is unaffected by this change.   The autolocker works fine.

   When the power to the MC is increased, these will have to be replaced or else the WFS will burn.

ELOG V3.1.3-